OpenStudy (anonymous):

Need help me prove: Thanks! sin (A - B) cos 2B + cos (A - B) sin 2B = sin (A + B)

OpenStudy (blurbendy):

use identities

OpenStudy (dumbcow):

note: (A-B) + 2B = A+B let x = A-B y = 2B using angle sum identity for sine: \[\sin(x+y) = \sin x \cos y + \cos x \sin y\] plug back in values for x,y and you have proved the identity

OpenStudy (anonymous):

sin (A - B) cos 2B + cos (A - B) sin 2B = sin (A + B) LHS= sin (A - B) cos 2B + cos (A - B) sin 2B [by using sin(A+B)=sinAcosB+cosAsinB formula] =sin(A-B+2B) =sin(A+B) =RHS Thus sin (A - B) cos 2B + cos (A - B) sin 2B = sin (A + B)

OpenStudy (anonymous):

Thanks!